fling aginst the wind,mr.bob's plane flew 450 miles from one airport to another in 1.5 hours. the return trip,with no change of wind or direction, took 1.25 hours. what was the speed of the wind?
Answer by checkley71(8403)

450/1.5=300MPH & 450/1.25=360MPH THEREFORE THE WIND IS (360-300)/2 OR 60/2=30MPH IS THE WIND SPEED